# Transportability of Prognostic Markers: Rethinking Common Practices through a Sufficient-Component-Cause Perspective

### v1

@@ Abstract
We modeled population-generating scenarios similar to those presented in Figure 1. Because in realistic settings all three causes are likely to vary, we explored this setup with more depth. We modeled scenarios where changes in the three causes had different strengths of correlation with each other (ranging from changes being uncorrelated to fully correlated). We also modeled a ‘maximum entropy’ scenario where the prevalence of the eight subgroups follows Dirichlet(1,1,1,1,1,1,1,1); this distribution has maximum information-theoretic randomness (entropy)27. For each scenario, we simulated random pairs of source and target populations according to a given population-generating mechanism. For each pair, we applied the following three transportation strategies: by accuracy (SE and SP), by predictive values (PPV and NPV), and via the proportional-odds assumption explained previously. Divergence metrics were averaged across 10,000 simulations for each scenario.
@@
Results are provided in Figure 2. This confirms that if the outcome prevalence variation is entirely due to $T$, PPV and NPV are fully transportable ($D_{KL}=0$). However, under other scenarios, transporting by predictive values resulted in substantial information loss. Transporting by accuracy reduced information loss in some scenarios, albeit in none was the loss zero, indicating that none of the modeled scenarios were compatible with stable SE and SP. Transporting by accuracy performed worse than transporting by predictive values in some scenarios, indicating that prevalence adjustment is not guaranteed to improve the transportability of a prognostic marker. The proportional-odds assumption was also reduced information loss in some scenarios, but increased it in others. For the scenarios where the three causes changed together, transportation by accuracy generally reduced information loss compared with transportation by predictive values (unless the correlation among causes were weak). In all those situations, transportation via he proportional-odds assumption performed better than the conventional transportability methods. As expected, in the scenarios where changes in causes were positively correlated.
@@
  • The base population is as follows: $P(T)=0.25$, $P(U)=0.75$, and $P(V)=0.5$. When a cause varies, it is given a logit-normal distribution with mean being equal to the base population’s point estimate, and standard deviaioon of 0.1.
  • No coorelation, weak, moderate, strong, and full correlation correspond to correlation coefficients of, respectively, 0, 0.25, 0.50, 0.75, amd 1 for logit-transformed probabilities.
@@ Discussion
We constructed a simple causal model under the SCC framework to study the transportability of prognostic markers that are partially informative about an outcome of interest but also generate false-positive and false-negative results. Several observations from this theoretical exploration deserve being highlighted. First, the common practice of transporting the marker by predictive values is questionable, as it relies on the strict assumption that the prevalences of causes other than those represented by the marker remain stable across populations. Further, we showed that the conventional prevalence adjustment method is the exact implementation of Bayes’ rule, thus changing the locus of transportability from predictive values to accuracy metrics (sensitivity and specificity). Nevertheless, no easily explainable or biologically plausible mechanism is likely to generate fully transportable accuracy metrics. In our explored scenarios regarding the distribution of causes across populations, transporting a prognostic marker by accuracy was not universally better than transporting by predictive values. However, in scenarios where causes were positively correlated, which is likely to be common, transporting by accuracy did reduce information loss compared with transportability by predictive values.
@@
These results are based on a selected set of population-generating mechanisms and a single set of core parameters for the SCC model ($P(T)=0.25$, $P(U)=0.75$, $P(V)=0.5$). Still, they should be sufficient to question some common assumptions and practices. These include the practice of advertising prognostic risk questions without emphasizing their dependence on prevalence, or considering sensitivity and specificity as intrinsic properties of markers. While these insights are from studying binary markers, the core findings can be extended to continuous markers and multi-variable risk equations. Transforming such markers ‘as is’ is equal to assuming that stratum-specific predictive values ($P(D|T=x)$) remain constant. This is equal to attributing variations in prevalence entirely to variations in $P(T)$ (aka case-mix). Consider a multi-variable risk score such as the QRISK for cardiovascular diseases28. Even though this model includes up to 22 predictors, is it plausible that other causal pathways to cardiovascular diseases that might vary across populations, including yet-to-be-discovered genetic risk factors, environmental exposures, subtle lifestyle choices, access to and quality of preventive care, and so on, are the same across populations? This assumption is needed for claiming that QRISK predictions, developed using primary care UK data, are transportable to other settings. On the other hand, conventional prevalence adjustment would indicate that stratum-specific likelihood ratio, i.e., $P(x|D)/P(x|\dot{D})$ would be transportable. This assumption will hold if the distribution of predicted risks among those who will experience the outcome and those who will not remains is same across diverse populations. This condition places very specific constraints on how the causes of diseases should vary across populations.
@@
We focused on prognostic markers given the recent debates on transportabiltiy of prognostic information and the merit of prevalence-adjustment to improve transportability. The application of SCC framework for diagnostic markers is also important and deserves its own airing. For such markers, the path of causality if from the disease to the test. As such, an equivalent setup to our reference setup for a diagnostic marker would be $T=(D\lor V)\land U$. The conventional method of transporting diagnostic markers is by accuracy (e.g., the Bayes’ rule, which combines accuracy metrics from the source population with disease prevalence from the new population). From this setup, it is immediately obvious that transportation by accuracy requires that universal and alternative causes $U$ and $V$ remain stable across populations. This assumption is similar to the assumption of transportability by predictive values for prognostic markers. This strong requirement can explain why these metrics tend to vary across populations29–31. Further details of how the SCC framework can inform transportability of diagnostic markers needs to be pursued in future work.
@@
How can these findings inform practice? If reliable, context-specific information is available on the distribution of specific predictors and their relationships with the outcome and with each other, methods based on causal graphs can be used to build transportable models or to design tailored transportation strategies11. In contrast, when only general information, such as outcome prevalence or test positivity rates are available, the SCC framework can be used to formulate tailored transportability algorithms that utilize such information. We question the contemporary practice of advertising risk prediction models for outcome risk as transportable. We believe risk equations should explicitly mention the outcome risk in their source population, and prevalence adjustment should be advocated as a necessary step before using the equation in a population with different outcome prevalence. Yet, while conventional prevalence adjustment may serve as a reasonable default, it cannot be assumed to consistently improve performance. Instead of fixed transportability rules applied to common performance metrics, the SCC framework offers a foundation for updating marker performance in ways that are more grounded in mechanistic interactions among potential causes. This results in new transportability algorithms based on assumptions on how underlying causes vary, rather than keeping performance metrics, which are emergent properties of such causes, constant. An example is the proportional-odds prevalence adjustment method discussed above. It is motivated by the notion that oftentimes differences in outcome prevalence are multi-factorial. An agnostic stance that that treats all causes as equally responsible is more defendable than blaming only one (as in transportation by predictive values) or assuming they move in unxplainable ways just to keep sensitivity and specificity constant (as in transportability by accuracy). Of course, variations in this method are also possible. Imagine, for example, that in addition to knowing target prevalence, we have knowledge of marker positivity rate ($P(T)$) in the new population (e.g., from an independent cross-sectional study). One can then directly adjust $P(T)$ to match this value, and then apply a common odds-ratio to other causes to match target prevalence.
@@
Appendix 1: Operationalizing the three types of transportation
Let {TP,FP,FN,TNTP,FP,FN,TN} be the elements of the two-by-two contingency table from the source population. We would like to transport this marker to a new setting where prevalence is $\pi$. Our goal is to construct a predicted contingency table for the target population, defined by $TP^{\prime},FP^{\prime},FN^{\prime},TN^{\prime}$}.
A. Transportation by accuracy:
@@
    Compute sensitivity and specificity from the source population
    $SE=TP/(TP+FN)$, $SP=TN/(FP+TN)$
@@
    Combine the SE and SP from the source population with prevalence from the target population
    $TP^{\prime}=\pi SE,\qquad FP^{\prime}=(1-\pi)(1-SP),\qquad FN^{\prime}=\pi(1-SE),\qquad TN^{\prime}=(1-\pi)SP$
B. Transportation by predictive values:
  1. 1.
    Compute positive and negative predictive values from the source population
    $PPV=TP/(TP+FP)$, $NPV=TN/(FN+TN)$
  2. 2.
    Compute $p$, the probability of positive marker results in the new population. Given that $\pi=pPPV+(1-p)(1-NPV)$, we have $p=(\pi-1+NPV)/(PPV+1-NPV)$.
  3. 3.
    Combine the PPV and NPV from the source population with test positivity from the target population
    $TP^{\prime}=pPPV,\qquad FP^{\prime}=p(1-PPV),\qquad FN^{\prime}=(1-p)(1-NPV),\qquad TN^{\prime}=(1-p)NPV$
C. Transportation by the proportional-odds assumption:
  1. 1.
    Map from {TP,FP,FN,TN} to {T,U,V}:
    $T=TP+FP,\qquad U=TP/(TP+FP),\qquad V=FN/((1-T)U)$
  2. 2.
    Find the odds-ratio $x$ such that when applied to $T$, $U$, $V$, results in updated probabilities $T^{\prime}$, $U^{\prime}$, $V^{\prime}$ that correspond to the desired prevalence. Note that prevalence is equal to $T^{\prime}(U^{\prime}+V^{\prime}-U^{\prime}V^{\prime})$.
@@ C. Transportation by the proportional-odds assumption:
    Recalculate the contingency table from $T^{\prime}$, $U^{\prime}$, $V^{\prime}$:
    $TP^{\prime}=T^{\prime}U^{\prime},\qquad FP^{\prime}=T^{\prime}(1-U^{\prime}),\qquad FN^{\prime}=(1-T^{\prime})U^{\prime}V^{\prime},\qquad TN^{\prime}=1-TP^{\prime}-FP^{\prime}-FN^{\prime}$

### v2

@@ Abstract
We modeled population-generating scenarios similar to those presented in Figure 1. Because in many realistic settings all three causes are likely to vary, we explored this setup with more depth. We modeled scenarios where changes in the three causes had different strengths of correlation with each other (ranging from changes being uncorrelated to fully correlated). We also modeled a ‘maximum entropy’ scenario where the three causes have independent uniform distributions; albeit unrealistic, this scenario represents maximum theoretical randomness across populations (therefore, marker performance is entirely untransportable)27. For each scenario, we simulated random pairs of source and target populations according to a given population-generating mechanism. When a cause was made variable, it was assumed that the logit-transformed probability of that cause had a standard Normal distribution (e.g., $\mbox{logit}(P_{T})\sim\mbox{Normal}(0,1)$ when where $P_{T}$ was variable across populations). For each pair, we applied the following three transportation strategies: by accuracy (SE and SP), by predictive values (PPV and NPV), and via the proportional-odds assumption explained previously. Divergence metrics were averaged across 10,000 simulations for each scenario.
@@
Results are provided in Figure 2. This confirms that if the outcome prevalence variation is entirely due to $T$, PPV and NPV are fully transportable ($D_{KL}=0$). However, under other scenarios, transportation by predictive values resulted in information loss, sometimes substantially. Transportation by accuracy reduced information loss in some scenarios, but increased in others. This indicates that prevalence adjustment is not guaranteed to improve the transportability of a prognostic marker. In none of the scenarios was the loss zero, indicating that none of the modeled scenarios were compatible with stable SE and SP - again, due to the complex requirements for the stability of these metrics. The proportional-odds assumption also reduced information loss in some scenarios, but increased it in others. For the scenarios where the three causes changed together, transportation by accuracy generally reduced information loss compared with transportation by predictive values. In such scenarios, transportation via the proportional-odds assumption performed better than both conventional transportability methods.
@@
  • Aside in the maximum entropy scenario (whose description provided in text), he variable component in each scenario had a standard logit-normal distribution e.g., $\mbox{logit}(T)\propto\mbox{Normal}(0,1)$ for scenarios where $T$ was variable.
  • No correlation, weak, moderate, strong, and full correlation correspond to correlation coefficients of, respectively, 0, 0.25, 0.50, 0.75, amd 1 for logit-transformed probabilities.
@@ Discussion
We constructed a simple causal model under the SCC framework to study the transportability of prognostic markers that are partially informative about an outcome of interest but also generate false-positive and false-negative results. Several observations from this theoretical exploration deserve being highlighted. First, the common practice of transporting the marker by predictive values is questionable, as it relies on the strict assumption that the prevalences of causes other than those represented by the marker remain stable across populations. Further, we showed that the conventional prevalence adjustment method is the exact implementation of Bayes’ rule, thus changing the locus of transportability from predictive values to accuracy metrics (sensitivity and specificity). Nevertheless, no easily explainable or biologically plausible mechanism is likely to generate fully transportable accuracy metrics. In our explored scenarios regarding the distribution of causes across populations, transporting a prognostic marker by accuracy was not universally better than transporting by predictive values. However, in scenarios where causes were positively correlated, which is likely to be common, transporting by accuracy did reduce information loss compared with transportability by predictive values. However, in the same scenarios, the SSC-based ‘proportional odds’ method of transportation performed better.
@@
Our simulations were related to a selected scenarios and particular ways the causes vary across populations. Still, they should be sufficient to question some common assumptions and practices. These include the practice of advertising prognostic risk questions without emphasizing their dependence on prevalence, or considering sensitivity and specificity as intrinsic properties of markers. While these insights are from studying binary markers, the core findings can be extended to continuous markers and multi-variable risk equations. Transforming such markers ‘as is’ is equal to assuming that stratum-specific predictive values ($P(D|T)$) remain constant. This is equal to attributing variations in prevalence entirely to variations in $P_{T}$ (aka case-mix). Consider a multi-variable risk score such as the QRISK for cardiovascular diseases28. Even though this model includes up to 22 predictors, is it plausible that other causal pathways to cardiovascular diseases that might vary across populations, including environmental exposures, subtle lifestyle choices, access to and quality of preventive care, and yet-to-be-discovered genetic risk factors, are the same across populations? This assumption is needed for claiming that QRISK predictions, developed using primary care UK data, are transportable to other settings. On the other hand, conventional prevalence adjustment would indicate that stratum-specific likelihood ratio, i.e., $P(x|D=1)/P(x|D=0)$ would be transportable. This assumption will hold if the distribution of predicted risks among those who will experience the outcome and those who will not remains the same across diverse populations. This condition places very specific constraints on how the causes of diseases should vary across populations.
@@
We focused on prognostic markers given the recent debates on transportability of prognostic information and the merit of prevalence-adjustment to improve transportability. The application of SCC framework for diagnostic markers is also important and deserves its own airing. For such markers, the path of causality if from the disease to the test. As such, an equivalent setup to our reference setup for a diagnostic marker would be $T=(D\lor V)\land U$. The conventional method of transporting diagnostic markers is by accuracy (e.g., the Bayes’ rule, which combines accuracy metrics from the source population with disease prevalence from the new population). From this setup, it is immediately obvious that transportation by accuracy requires that universal and alternative causes $U$ and $V$ remain stable across populations. This assumption is similar to the assumption of transportability by predictive values for prognostic markers. This strong requirement can explain why sensitivity and specificity tend to vary across populations29–31. How the SCC framework can inform transportability of diagnostic markers needs to be pursued in future work.
@@
How can these findings inform practice? If reliable, context-specific information is available on the distribution of specific predictors and their relationships with the outcome and with each other, methods based on causal graphs can be used to build transportable models or to design tailored transportation strategies11. In contrast, when only general information, such as outcome prevalence or test positivity rates, are available, the SCC framework can be used to formulate transportability algorithms that utilize such broad information. We question the contemporary practice of advertising risk prediction models for outcome risk as transportable. We believe risk equations should explicitly mention the outcome risk in their source population, and adjustment for prevalence should be advocated as a necessary step before using the equation in a population with different outcome prevalence. Yet, while conventional prevalence adjustment may serve as a reasonable choice, it cannot be assumed to consistently improve performance. Instead of fixed transportability rules, the SCC framework offers a foundation for updating marker performance in ways that are more grounded in mechanistic interactions among potential causes. This results in new transportability algorithms based on assumptions on how underlying causes vary, rather than keeping performance metrics, which are emergent properties of such causes, constant. An example is the proportional-odds prevalence adjustment method discussed above. It is motivated by the notion that oftentimes differences in outcome prevalence are multi-factorial. An agnostic stance that treats all causes as equally responsible is more defendable than blaming only one (as in transportation by predictive values) or assuming they move in unexplainable ways just to keep sensitivity and specificity constant (as in transportability by accuracy). Interestingly, in our numerical experiments, whenever conventional prevalence-adjustment performed better than transportability by predictive values, the SCC-based proportional-odds method performed better than both, indicating it might be a better choice. Of course, variations in this method are also possible. Imagine, for example, that in addition to knowing target prevalence, we have knowledge of marker positivity rate ($P_{T}$) in the new population (e.g., from an independent cross-sectional study). One can then directly adjust $P_{T}$ to match this value, and then apply a common odds-ratio to other causes to match target prevalence.
@@
Appendix: Transportation by the proportional-odds assumption
Let {TP,FP,FN,TNTP,FP,FN,TN} be the elements of the two-by-two contingency table. We would like to transport this marker to a new setting where prevalence is $\pi^{*}$. Our goal is to construct a predicted contingency table for the target population, defined by $TP^{*},FP^{*},FN^{*},TN^{*}$}.
@@
    Map from {TP,FP,FN,TNTP,FP,FN,TN} to $P_{T},P_{U},P_{V}$} (see text)
@@
    Find the odds-ratio $x$ such that when applied to $P_{T}$, $P_{U}$, $P_{V}$, results in updated probabilities $P_{T}^{*}$, $P_{U}^{*}$, $P_{V}^{*}$ that correspond to the desired prevalence. Note that prevalence is equal to $P_{U}^{*}(P_{T}^{*}+P_{V}^{*}-P_{T}^{*}P_{V}^{*})$.
@@
    Map the resulting $P_{T}^{*}$, $P_{U}^{*}$, $P_{V}^{*}$} to $TP^{*},FP^{*},FN^{*},TN^{*}$} (see text)